A mean of 3 out of 5 students request for tomato sauce with their meal in a canteen. A random sample of 8 students is selected, what is the probability that less than 4 students request for tomato sauce with their meal?



1
Expert's answer
2022-02-14T16:51:49-0500

Let "X=" the number of students request for tomato sauce with their meal in a cantee: "X\sim Bin(n, p)."

Given "n=8, p=3\/5=0.6, q=1-p=1-0.6=0.4."

i)


"P(X<4)=P(X=0)+P(X=1)"

"+P(X=2)+P(X=3)"

"=\dbinom{8}{0}(0.6)^0(0.4)^{8-0}+\dbinom{8}{1}(0.6)^1(0.4)^{8-1}"

"+\dbinom{8}{2}(0.6)^2(0.4)^{8-2}+\dbinom{8}{3}(0.6)^3(0.4)^{8-3}"

"=0.00065536+0.00786432"

"+0.04128768+0.12386304"

"=0.1736704"